OpenStudy (anonymous):

help! verify the identity: (sec(theta) + tan(theta)) (1 - sin(theta)) = cos(theta)

OpenStudy (marcelie):

you can distribute

OpenStudy (unklerhaukus):

LHS = (sec + tan) (1 - sin) = (1/cos + sin/cos)(1-sin) = 1/cos + sin/cos - sin/cos - sin^2/cos = =

OpenStudy (unklerhaukus):

simplify, and then use sin^2+cos^2 = 1